# On products of abelian skew braces

### v2

@@ 1 Introduction
Note that a weaker notion of factorised subbrace is presented in [9, Definition 2.17].
@@ 3 Proof of Theorem A
Since $[E_{1},K]=[E_{1},L_{1}L_{2}]=[E_{1},L_{2}]$ and $[E,L_{2}]=[E_{1}E_{2},L_{2}]=[E_{1},L_{2}]$, $K$ and $E$ normalises $[E_{1},L_{2}]$ by [8, Kapitel III, Hilfssatz 1.6]. Hence $[E_{1},L_{2}]$ is normal in $KE=G$. Analogously, $[E_{2},L_{1}]$ is normal in $G$.
@@
Skew brace version of the above result provides a slightly stronger version of [11, Proposition 3.3].
@@
Let us denote by $T_{i}=L_{i}H\cap L_{i}E$, $i=1,2$. Then $L_{i}E_{i}=L_{i}(L_{i}H\cap E)=T_{i}$, $i=1,2$. By Statement 3, $E_{i}$ is an abelian subgroup of $G$ and $[E_{i},L_{i}]\leq[T_{i},T_{i}]=1$, $i=1,2$. By Statement 2, $E_{2}$ normalizes $L_{1}$. Therefore, $T_{2}T_{1}=L_{2}E_{2}L_{1}E_{1}=L_{2}L_{1}E_{2}E_{1}=KE=G$ and $G$ is the product of two abelian groups. By Itô’s theorem ([8, Kapitel VI, Satz 4.4]), $G^{\prime}$ is abelian. In particular, $K^{\prime}[E,K]$ is abelian.
@@
Moreover, since $K=L_{1}L_{2}$ and $L_{1},L_{2}$ are abelian, it follows that $K^{\prime}=[L_{1},L_{2}]$. As $[[E_{1},L_{1}],L_{2}]=1$ ($[E_{1},L_{1}]=1$) and $[[L_{2},E_{1}],L_{1}]=1$ ($[L_{2},E_{1}]\leq L_{1}$), it follows that $[[L_{1},L_{2}],E_{1}]=1$ by the three subgroups lemma ([8, Kapitel III, Hilfssatz 1.10]). Hence $[E_{1},K^{\prime}]=1$. Since $G^{\prime}$ is abelian, $[E^{\prime}E_{1},K^{\prime}[E,K]]=[E_{1},K^{\prime}]=1$ (note that $[[E,K],E_{1}]\leq[L_{1},E_{1}]=1$). Consequently, $[K^{\prime}[E,K]H\cap E,K^{\prime}[E,K]]=1$, as required. ∎
@@ 5 Proof of Theorem C
Arguing as in Proposition 3.5, we have that $G=T_{1}T_{2}$ and $G^{\prime}$ is abelian by Itô’s theorem ([8, Kapitel VI, Satz 4.4]). Since $G$ is non-trivial, $K$ is non-trivial, therefore, $L_{1}$ or $L_{2}$ is non-trivial. If $K^{\prime}[E,K]=1$, every subgroup of $K$ is normal in $KE=G$. Hence $L_{1}$ and $L_{2}$ are normal subgroups of $G$, and the result is proved. Hence we may suppose that $K^{\prime}[E,K]\neq 1$.

### v3

@@ 1 Introduction
Note that a weaker notion of factorised subbrace is presented in [10, Definition 2.17].
@@ 3 Proof of Theorem A
The next result appears in [7, Propositions 2.1 and 2.2]. We give a slightly different proof for the sake of completeness.
@@
Since $[E_{1},K]=[E_{1},L_{1}L_{2}]=[E_{1},L_{2}]$ and $[E,L_{2}]=[E_{1}E_{2},L_{2}]=[E_{1},L_{2}]$, $K$ and $E$ normalises $[E_{1},L_{2}]$ by [9, Kapitel III, Hilfssatz 1.6]. Hence $[E_{1},L_{2}]$ is normal in $KE=G$. Analogously, $[E_{2},L_{1}]$ is normal in $G$.
@@
Skew brace version of the above result provides a slightly stronger version of [12, Proposition 3.3].
@@
Let us denote by $T_{i}=L_{i}H\cap L_{i}E$, $i=1,2$. Then $L_{i}E_{i}=L_{i}(L_{i}H\cap E)=T_{i}$, $i=1,2$. By Statement 3, $E_{i}$ is an abelian subgroup of $G$ and $[E_{i},L_{i}]\leq[T_{i},T_{i}]=1$, $i=1,2$. By Statement 2, $E_{2}$ normalizes $L_{1}$. Therefore, $T_{2}T_{1}=L_{2}E_{2}L_{1}E_{1}=L_{2}L_{1}E_{2}E_{1}=KE=G$ and $G$ is the product of two abelian groups. By Itô’s theorem ([9, Kapitel VI, Satz 4.4]), $G^{\prime}$ is abelian. In particular, $K^{\prime}[E,K]$ is abelian.
@@
Moreover, since $K=L_{1}L_{2}$ and $L_{1},L_{2}$ are abelian, it follows that $K^{\prime}=[L_{1},L_{2}]$. As $[[E_{1},L_{1}],L_{2}]=1$ ($[E_{1},L_{1}]=1$) and $[[L_{2},E_{1}],L_{1}]=1$ ($[L_{2},E_{1}]\leq L_{1}$), it follows that $[[L_{1},L_{2}],E_{1}]=1$ by the three subgroups lemma ([9, Kapitel III, Hilfssatz 1.10]). Hence $[E_{1},K^{\prime}]=1$. Since $G^{\prime}$ is abelian, $[E^{\prime}E_{1},K^{\prime}[E,K]]=[E_{1},K^{\prime}]=1$ (note that $[[E,K],E_{1}]\leq[L_{1},E_{1}]=1$). Consequently, $[K^{\prime}[E,K]H\cap E,K^{\prime}[E,K]]=1$, as required. ∎
@@ 5 Proof of Theorem C
Arguing as in Proposition 3.5, we have that $G=T_{1}T_{2}$ and $G^{\prime}$ is abelian by Itô’s theorem ([9, Kapitel VI, Satz 4.4]). Since $G$ is non-trivial, $K$ is non-trivial, therefore, $L_{1}$ or $L_{2}$ is non-trivial. If $K^{\prime}[E,K]=1$, every subgroup of $K$ is normal in $KE=G$. Hence $L_{1}$ and $L_{2}$ are normal subgroups of $G$, and the result is proved. Hence we may suppose that $K^{\prime}[E,K]\neq 1$.